remi2019:

A line passes through (3, 7) and (6, 9). Which equation best represents the line?

remi2019:

plz help

AnimeLover4363:

y = 3/2 x 5 y = 2/3 x 5 y = 3x 2 y = 2/3 x 2 i think this is it

XxAngelGirlxX:

1 attachment
AnimeLover4363:

well thats just the simple answer lol